# Representing local protein environments with atomistic foundation models

### v2

@@ 3 Representing protein local environments with AFMs
Notation. Denote a residue in a protein as ${a}$, and the all-atom representation of the protein structure as $\mathcal{X}=\{(\mathbf{x}_{1},z_{1})\dots(\mathbf{x}_{m},z_{m})\}$, where $\mathbf{x}_{j}\in\mathbb{R}^{3}$ and $z_{j}$ are the location and atomic number of $j^{th}$ atom in $\mathcal{X}$, respectively. Let $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ be the AFM with parameters $\theta$, and $\mathcal{Y}=\{(\mathbf{y}_{1},z_{1}),\ldots,(\mathbf{y}_{m},z_{m})\}$ the atom-wise feature representations of the AFM, where $\mathbf{y}\in\mathbb{R}^{d}$. The local environment of a focus residue ${a}$ is denoted by $\mathcal{X}_{{a}}\subseteq\mathcal{X}$. The subset of atoms in residue ${a}$ used for predicting biochemical properties is denoted as $\mathcal{A}_{a}$. Given a subset of atom indices $\mathcal{A}_{a}$, we denote by $\mathcal{Y}_{\mathcal{A}_{a}}=\{\,(\mathbf{y}_{i},z_{i})\mid i\in\mathcal{A}_{a}\}\subseteq\mathcal{Y}$ the restriction of $\mathcal{Y}$ to those atoms.
@@
We seek a representation of a local protein environment to be (i) sensitive to local changes, (ii) insensitive to global variations, (iii) fast to compute, (iv) canonical, enabling direct comparison across diverse environments, and (v) effective and generalizable to unseen environments. Encoding full proteins with several thousand atoms with AFMs is computationally inefficient and redundant since we are solely interested in locally sensitive features. To construct such representations, we propose to construct a local environment around a focus residue to be encoded later by the AFM.
@@
Creating local environments. To balance computational efficiency while retaining local structural context, given a focus residue $a$, we construct the environment $\mathcal{X}_{a}$ as the union of all residues whose atoms are at most $5$Å away from the atoms of ${a}$. The procedure is described in Algorithm  1.
@@
AFM representations. Given an environment $\mathcal{X}_{a}$, AFMs produce atom-level features over the layers of the network. These learned atomistic features are contextual, influenced not only by the atom itself but also by the other atoms in the environment due to the message-passing operations within AFMs. Given $\mathcal{X}_{a}$, atom-wise feature representations are extracted from the final layer of the AFM. Each AFM produces representations of shape $N\times d$, where $N=|\mathcal{X}_{a}|$ and $d$ is the dimension of the representation. In order to construct representations that are canonical, i.e., comparable across residues, the representations corresponding to atoms in $\mathcal{A}$ are retained, yielding contextualized representations of their local chemical environments per residue. Given a protein structure, this process is performed for each residue in the sequence, and residue-wise features are curated. A visual depiction of this procedure is presented in Fig. 1.
4 AFM representations capture local protein structure and chemistry
In what follows, we investigate how AFM representations of local environments can be leveraged to predict a range of structural and chemical properties.
@@ 4 AFM representations capture local protein structure and chemistry
Amino acid identity & secondary structure. A protein’s primary structure is a linear sequence of amino acids connected by covalent peptide bonds. The twenty standard proteinogenic amino acids kawashima2000aaindex share a backbone of four heavy atoms: the $\alpha$-carbon (CA), amide nitrogen (N), carbonyl carbon (C), and carbonyl oxygen (O), with variable side chains (R-group). A protein’s secondary structure is a local motif of backbone conformation, stabilized by hydrogen bonds between atoms of non-adjacent amino acids kabsch1983dictionary , with the most common types being $\alpha$-helices (H) and $\beta$-strands (E).
@@
To examine the information encoded by the AFM representations, the atomistic features from MACE were projected into a two dimensional space using Uniform Manifold Approximation and Projection (UMAP) mcinnes2018umap . The resulting embeddings, shown in Fig. A.8, are annotated with amino acid labels, secondary structure, and dihedral angles. Notably, features corresponding to $\alpha$-helices and $\beta$-strands form distinct clusters in the UMAP space. A similar pattern of separation is observed when embeddings are annotated by amino acid identity and backbone dihedral angles, indicating that chemically and structurally distinct features are well-represented in the feature space.
@@
To quantitatively assess this capability, we train classifiers to predict the secondary structures and the amino acid identities. Further details on the loss function, target label space, and the definition of $\mathcal{A}_{a}$ can be found in Appendix F.1, while a comprehensive description of the model architecture is provided in Appendix F.3. Classifier accuracy was evaluated using precision, recall, and F1 scores. Results for secondary structure prediction are summarized in Table B.1, while amino acid classification results are presented in Table B.2. For secondary structure, models trained with MACE and Egret features consistently achieve superior prediction on average. For amino acid prediction, models trained using Egret features consistently outperform those trained with alternative representations.
Acid dissociation constants (pKa). pKa quantifies the tendency of a titratable group (such as the carboxylic acid group of an Asp side chain) to donate a proton, serving as a measure of its acidity. When the solvent pH is lower than pKa, the molecule will be primarily in its protonated form, while being increasingly deprotonated at pH higher than pKa. The pKa value is the pH value at which the group is half protonated over the ensemble of molecules. Although each amino acid has a canonical pKa value in isolation, its actual value within a protein can deviate due to variations in the local chemical environment. External factors such as (i) the presence of nearby charged residues, polar atoms, or functional groups, and (ii) limited solvent accessibility—particularly in buried regions—can influence protonation behavior. This makes pKa prediction an ideal test case for evaluating whether atomistic features can encode meaningful local chemical information.
@@
For quantitative assessment, a regression model was trained on amino acids with ionizable, proton-transferring side chains to predict their pKa values. Ground-truth pKa values were estimated using a Poisson–Boltzmann solver reis2020pypka . The evaluation focused on glutamic acid (GLU) and aspartic acid (ASP), which tend to deprotonate, and lysine (LYS) and histidine (HIS), which tend to protonate. Additional training details are provided in Appendix F. As shown in Table B.3, models trained with AIMNet features outperform those using alternative features because it is trained to predict multiple molecular properties simultaneously. All AFM-based regressors outperform conventional baselines such as PropKA olsson2011propka3 and pKa-ANI gokcan2022prediction .
Figure 3: Chemical shift prediction accuracy of CA (left) and N (right) atoms stratified by the KDE-estimated likelihood of the corresponding MACE descriptors. Depicted are the median and $25\%$-$50\%$ confidence intervals. Higher-likelihood environments correspond to lower prediction error and can be used as an uncertainty measure.
5 Computing likelihoods and measuring similarity of protein environments
@@
The AFM embedding space encodes local biomolecular structure, capturing structural and chemical context. We leverage the richness of AFM embedding spaces to define a distribution over biomolecular environments. Given a set of reference environments $\mathcal{E}_{\text{ref}}=\{\mathcal{X}_{1},\ldots,\mathcal{X}_{n}\}$ and a reference atom set $\mathcal{A}$, we define the likelihood of an environment $\mathcal{X}_{a}$ at a focus residue $a$ as
@@
where $\sigma$ refers to the bandwidth, and it controls the influence of each reference environment on $\mathcal{X}_{a}$. This is equivalent to performing kernel-density estimation in the AFM embedding space, with a radial basis function kernel. Intuitively, the likelihood measures how typical $\mathcal{X}_{a}$ is among the reference environments $\mathcal{E}_{\text{ref}}$. Subsequently, the conditional likelihoods can be defined similarly to Eq. 1 by curating an appropriate set of reference environments that satisfy the chosen conditioning.
Measuring distribution shifts. To evaluate the quality of unconditional likelihood estimation and its ability to detect subtle distribution shifts in the local structure, we curated a set of reference environments randomly sampling $10,000$ environments from $1,100$ protein structures that were relaxed using the Amber99 force field hornak2006comparison . We then measured the likelihoods of each environment sampled from a test set of $225$ proteins, before and after performing Amber99 relaxation. The results shown in Fig. 5 demonstrate that the likelihood function is sensitive to subtle conformational changes. Relaxed structures consistently receive higher likelihoods, and the distribution of the paired differences captures fine-grained structural variations. This makes the approach well-suited for detecting out-of-distribution conformations and assessing local structural quality.
Secondary-structure conditioned likelihoods. To further evaluate the sensitivity of conditional likelihoods, we randomly sampled $1000$ environments from secondary structures annotated as $\alpha$-helix (H), $\beta$-strand (E), and turn (T) by DSSP, and constructed conditional likelihoods $p(\mathcal{X}|{\text{H}})$, $p(\mathcal{X}|{\text{E}})$, and $p(\mathcal{X}|\text{T})$, respectively. Fig. A.9 depicts the projections of conditional likelihood triplets $\left(\log p(\mathcal{X}_{i}|\mathrm{E}),\log p(\mathcal{X}_{i}|\mathrm{H}),\log p(\mathcal{X}_{i}|\mathrm{T})\right)$ of $8,163$ environments from the test set. The results highlight a clear separation by true secondary structure of the environment. The one-dimensional marginal distributions further emphasize the distinct statistical profiles of each structural class.
We further direct the reader to Fig. A.12 and Appendix G.1 for a detailed analysis of the use of AFM embeddings to evaluate chemical environment similarity.
@@
6 Physics-grounded chemical shift prediction for proteins
@@
Background and prior art. We refer an uninitiated reader to Appendix C for a short background on chemical shifts and biomolecular NMR. Several computational methods have been developed to predict NMR chemical shifts from molecular structures (shen2008sparta, ; neal2003shiftx, ; han2011shiftx2, ; Li2012ppm, ; Li2015ppmone, ; kohlhoff2009camshift, ; meiler2003proshift, ; moon2007shifts, ), with recent approaches increasingly relying on machine learning. The current state-of-the-art, UCBShift (ucbshift2023, ), combines sequence and structure alignment with a random forest model, using reference proteins and structural descriptors to guide predictions. Its latest iteration, UCBShift 2.0 (ptaszek2024ucbshift2, ), extends this framework to include side-chain atoms alongside backbone predictions. Despite their effectiveness, these methods depend on reference-based similarity measures, which limits their generalizability.
@@
AFM-based shift predictor. Chemical shift prediction is formulated as a downstream task over AFM embeddings. A lightweight graph neural network takes AFM representations as input and predicts the chemical shifts of a target atom. Separate models are trained for the N, CA, C, H, and HA backbone atoms, as well as the CB, CG, CD, CD2, CG2, and CE side-chain atoms. Dataset curation, model architecture, and training details are provided in Appendix D,  F.3.
@@
As evident from Fig. 4 , the proposed predictor outperforms UCBShift2 for both backbone and side-chain heavy atoms with the exception of the hydrogen HA. A comprehensive quantitative comparison of chemical shift prediction accuracy across different AFM variants is provided in Table. B.4. We observe that MACE-based models consistently achieve the best performance.
@@
To further probe the shift predictor and evaluate whether it correctly captures the local structural influence on the shift, we carefully design the following three case studies.
@@
Effects of ring currents. An external magnetic field induces a flow of delocalized $\pi$ electrons in an aromatic ring, which in turn produces its own magnetic field. Because chemical shifts depend on the net magnetic field produced by the local structural environment, they are strongly influenced by such ring currents. As the phenylalanine ring has a C2 symmetry, states separated by rotation of the ring by $180^{\circ}$ are indistinguishable. To probe whether our shift predictor captures ring-current effects accurately, we synthetically modified the side chain of a Phe (residue $50$ in PDB ID: 1ZV6) by rotating its aromatic ring ring about the $\chi_{2}$ dihedral angle from $-180^{\circ}$ to $180^{\circ}$. Fig. 2 depicts the studied environments and the predicted chemical shifts. Our AFM-based shift predictor shows the expected $180^{\circ}$ periodicity haigh1979ring in the chemical shifts of backbone CA atoms in the vicinity of the ring, with the ring current influence decaying smoothly as the distance from the ring increases and becoming negligible after  $7$Å. In contrast, UCBShift extends this influence beyond the expected range and fails to reproduce the smoothness and periodicity expected in theory.
Alternative conformations. We evaluate our shift predictor over a 100 ns MD simulation of the protein 4OLE:B from rosenberg2024seeingdouble , which features two stable conformers: conformer A containing a helix and conformer B with a linearly structured loop (Fig. A.11). Our method captures distinct secondary chemical shift distributions for lysine $63$ CA and CB atoms. The helical conformer exhibits the expected increase in CA secondary shift and decrease in CB secondary shift.
@@
Uncertainty estimation. We first build the unconditional likelihood model $p(\mathcal{X})$ for biomolecular environments on the training set used for shift prediction. We then evaluate this likelihood for every test environment and record the corresponding chemical shift prediction error. Fig. 3 reports the distribution of errors, stratified by likelihood. Environments with lower likelihoods consistently exhibit larger prediction errors, indicating that the likelihood can serve as a practical measure of confidence in chemical shift prediction tasks. To the best of our knowledge, ours is the first chemical shift predictor to supply such a confidence score.
@@
We summarize by concluding that AFM-based shift prediction not only delivers better accuracy, but is also physics-grounded and allows us to provide confidence estimates for each prediction.
Figure 5: MACE-based likelihood captures structurally subtle distribution shifts in protein environments. Left panel: the histogram of likelihoods of CA MACE representations of $26000$ protein environments curated from $225$ protein structures in the test set, before (red) and after (blue) relaxing the protein structures with an Amber99 force-field hornak2006comparison . Right panel: the relative likelihood of the same environment before and after relaxation.
7 Interpreting AFM representations

### v3

@@ 3 Representing protein local environments with MLFFs
Notation. We denote a residue in a protein as a{a}, and the all-atom representation of the protein structure as $\mathcal{X}=\{(\mathbf{x}_{1},z_{1})\dots(\mathbf{x}_{m},z_{m})\}$, where $\mathbf{x}_{j}\in\mathbb{R}^{3}$ and $z_{j}$ are the location and atomic number of $j^{th}$ atom in $\mathcal{X}$, respectively. Let $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ be the MLFF with parameters $\theta$, and $\mathcal{Y}=\{(\mathbf{y}_{1}),\ldots,(\mathbf{y}_{m})\}$ the atom-wise embeddings of the MLFF, where $\mathbf{y}\in\mathbb{R}^{d}$. The local environment of a focus residue a{a} is denoted by $\mathcal{X}_{{a}}\subseteq\mathcal{X}$. The subset of atoms in residue a{a} used for predicting biochemical properties is denoted as $\mathcal{A}_{a}$. Given a subset of atom indices $\mathcal{A}_{a}$, we denote by $\mathcal{Y}_{\mathcal{A}_{a}}=\{\,(\mathbf{y}_{i})\mid i\in\mathcal{A}_{a}\}\subseteq\mathcal{Y}$ the restriction of $\mathcal{Y}$ to those atoms.
@@
We seek a representation of a local protein environment to be (i) sensitive to local changes, (ii) insensitive to global variations, (iii) fast to compute, (iv) canonical, enabling direct comparison across diverse environments, and (v) effective and generalizable to unseen environments. Encoding full proteins with several thousand atoms with MLFFs is computationally inefficient and redundant since we are solely interested in locally sensitive features. To construct such representations, we propose to construct a local environment around a focus residue to be encoded later by the MLFF.
@@
Constructing local environments. To balance computational efficiency while retaining local structural context, given a focus residue $a$, we construct the environment $\mathcal{X}_{a}$ as the union of all residues whose atoms are at most 55Å away from the atoms of a{a}. The procedure is described in App. Alg. 1, and an ablation over the choice of radius is provided in Appendix F.
@@
MLFF representations. Given an environment $\mathcal{X}_{a}$, MLFFs produce atom-level features over the layers of the network. These learned atomistic features are contextual, influenced not only by the atom alone but also by the other atoms in the environment due to the message-passing operations within MLFFs. Given $\mathcal{X}_{a}$, atom-wise feature representations are extracted from the final layer of the MLFF. [^2] See Appendix F.2 for an ablation study comparing representations extracted from different MLFF layers. Each MLFF produces representations of shape $N\times d$, where $N=|\mathcal{X}_{a}|$ and $d$ is the dimension of the representation. To obtain canonical representations that are comparable across residues, we retain only $\mathcal{Y}_{\mathcal{A}_{a}}$, the representations corresponding to atoms in $\mathcal{A}$. Given a protein structure, this process is applied to each residue in the sequence to obtain residue-wise features. A visual depiction of this procedure is presented in Fig. 1.
Data. To ensure a consistent dataset across all experiments, we sourced data from RefDB, a curated subset of the Biomolecular Magnetic Resonance Databank (BMRB). The final dataset includes $1048$ BMRB entries, filtered for sequence redundancy and split into $823$ training and $225$ test proteins. From these structures, a total of 165165K local protein environments were extracted using a 55Å radius threshold based on Hausdorff distance. We use AlphaFold2 (Jumper et al., 2021) predicted 3D structures for all sequences, followed by hydrogen addition and Amber99 force-field relaxation (Hornak et al., 2006). We refer to Sec.  E for a detailed justification of our choice, and for further details on data curation.
Baselines To evaluate the utility of MLFF embeddings, we compare the performance of MLFF embeddings + GCN against different embedding baselines: (i) sequence-model embeddings (ESM3 (Hayes et al., 2025)) and sequence models with enhanced structural representations (Ouyang-Zhang et al., 2025) + GCN, (ii) ESMFold (Lin et al., 2023) structure-model embeddings + GCN, (iii) hand-crafted statistics-based descriptors (LOCO-HD) (Fazekas et al., 2024) + GCN, and (iv) simple learned embeddings + GCN. In the learned-embedding baseline, the GCN is trained end-to-end from atom types and 3D Cartesian coordinates from AlphaFold2 Jumper et al. (2021). The coordinates are first mapped to a sinusoidal positional encoding, which is then concatenated with the embedded atom-type identifiers to produce the node features for the GCN.
4 MLFF representations capture local protein structure and chemistry
In what follows, we first present a zero-shot analysis of MLFF representations to assess whether they inherently capture biochemically meaningful information without task-specific training. We then apply transfer learning by training neural networks on top of the frozen embeddings to predict structural and chemical properties. We first consider amino acid identity and secondary structure prediction as motivating examples, and then progress to more challenging downstream regression tasks such as pKa and NMR chemical shift prediction.
@@ 4 MLFF representations capture local protein structure and chemistry
Amino acid identity & Secondary structure. A protein’s primary structure is its linear sequence of amino acids linked by covalent peptide bonds and implicitly encode the 3D protein structure. Its secondary structure is a local motif of backbone conformation, stabilized by hydrogen bonds between non-adjacent residues (Kabsch & Sander, 1983), with the most common motifs being $\alpha$-helices (H) and $\beta$-sheets (E).
@@
To examine the information encoded by the MLFF representations in a zero-shot setting, the atomistic features from MACE were projected into a two dimensional space using Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018). The resulting embeddings, shown in Fig. 2, are annotated with amino acid labels, secondary structure, and dihedral angles. Notably, features corresponding to $\alpha$-helices and $\beta$-sheets form distinct clusters in the UMAP space. A similar pattern of separation is observed when embeddings are annotated by amino acid identity and backbone dihedral angles, indicating that chemically and structurally distinct features are well-represented in the feature space.
@@
To quantitatively assess this capability, we perform lightweight transfer learning by training classifiers to predict the secondary structures and the amino acid identities from the frozen embeddings. Further details on the loss function, target label space, and the definition of $\mathcal{A}_{a}$ can be found in Appendix G.1, while a comprehensive description of the model architecture is provided in Appendix G.3. Classifier accuracy was evaluated using precision, recall, and F1 scores. Results for secondary structure prediction are summarized in Table B.3, while amino acid classification results are presented in Table B.4. For secondary structure, models trained with MACE and Egret features consistently achieve superior prediction on average. For amino acid prediction, models trained using Egret features consistently outperform those trained with alternative representations.
Figure 2: MACE embedding space reveals meaningful structural and chemical information. Depicted are two-dimensional UMAP coordinates of $165,913$ protein environments from $1327$ non-redundant chains predicted by AlphaFold2 (Jumper et al., 2021), labeled left-to-right, top-to-bottom according to the DSSP secondary structure class (Table B.12), amino acid chemical identity, the pair of backbone dihedral angles $(\phi,\psi)$, and CA secondary chemical shift (relative to a random coil).
Acid dissociation constants (pKa). pKa quantifies the proton-donation propensity of a titratable group, i.e., its acidity. It is the pH at which the group is half protonated across the ensemble. In proteins, pKa values often deviate from their canonical solution values due to local electrostatics and solvent accessibility. These context-dependent shifts make pKa prediction a strong test of whether atomistic representations capture local chemical environments (see App. D for biochemical background).
@@
For quantitative assessment, a regression model was trained on amino acids with ionizable, proton-transferring side chains to predict their pKa values. Ground-truth pKa values were estimated using a Poisson-Boltzmann solver (Reis et al., 2020). The evaluation focused on glutamic acid (GLU) and aspartic acid (ASP), which tend to deprotonate, and lysine (LYS) and histidine (HIS), which tend to protonate. As shown in Table 1, models trained with AIMNet features achieve the best overall performance among all feature types because it is trained to predict multiple molecular properties simultaneously. For additional comparisons with PropKA (Olsson et al., 2011) and pKa-ANI (Gokcan & Isayev, 2022) on experimental pKa values and PDB structures, as well as training details, see Appendix G.
5 Likelihoods and similarity of local protein environments
@@
The MLFF embedding space encodes local biomolecular structure, capturing structural and chemical context. We leverage the richness of MLFF embedding spaces to define a distribution over biomolecular environments. Given a set of reference environments $\mathcal{E}_{\text{ref}}=\{\mathcal{X}_{1},\ldots,\mathcal{X}_{n}\}$ and a reference atom set $\mathcal{A}$, we define the likelihood of an environment $\mathcal{X}_{a}$ at a focus residue $a$ as
@@
where $\sigma$ refers to the bandwidth, and it controls the influence of each reference environment on $\mathcal{X}_{a}$. This is equivalent to performing kernel-density estimation in the MLFF embedding space, with a radial basis function kernel. Intuitively, the likelihood measures how typical $\mathcal{X}_{a}$ is among the reference environments $\mathcal{E}_{\text{ref}}$. Subsequently, the conditional likelihoods can be defined similarly to Eq. 1 by curating an appropriate set of reference environments that satisfy the chosen conditioning.
Capturing distribution shifts. To evaluate the quality of unconditional likelihood estimation and its ability to detect subtle distribution shifts in the local structure, we curated a set of reference environments randomly sampling $10,000$ environments from $1,100$ protein structures that were relaxed using the Amber99 force field (Hornak et al., 2006). We then measured the likelihoods of each environment sampled from a test set of $225$ proteins, before and after performing Amber99 relaxation. The results shown in Fig. A.7 demonstrate that the likelihood function is sensitive to subtle conformational changes. Relaxed structures consistently receive higher likelihoods, and the distribution of the paired differences captures fine-grained structural variations. This makes the approach well-suited for detecting out-of-distribution conformations and assessing local structural quality.
Secondary structure conditioned likelihoods. To further evaluate the sensitivity of conditional likelihoods, we randomly sampled $1000$ environments from secondary structures annotated as $\alpha$-helix (H), $\beta$-strand (E), and turn (T) by DSSP, and constructed conditional likelihoods $p(\mathcal{X}|{\text{H}})$, $p(\mathcal{X}|{\text{E}})$, and $p(\mathcal{X}|\text{T})$, respectively. Fig. A.11 depicts the projections of conditional likelihood triplets $\left(\log p(\mathcal{X}_{i}|\mathrm{E}),\log p(\mathcal{X}_{i}|\mathrm{H}),\log p(\mathcal{X}_{i}|\mathrm{T})\right)$ of $8,163$ environments from the test set. The results highlight a clear separation by true secondary structure of the environment. The one-dimensional marginal distributions further emphasize the distinct statistical profiles of each structural class.
We further direct the reader to Fig. A.13 and Appendix H.2 for a detailed analysis of the use of MLFF embeddings to evaluate chemical environment similarity.
@@
6 Physics-grounded and uncertainty-aware chemical shift prediction for proteins
@@
Background and prior art. We refer an uninitiated reader to Appendix D for a short background on chemical shifts and biomolecular NMR. Several computational methods have been developed to predict NMR chemical shifts from molecular structures (Shen & Bax, 2008; Neal et al., 2003; Han et al., 2011; Li & Brüschweiler, 2012; 2015; Kohlhoff et al., 2009; Meiler, 2003; Moon & Case, 2007), with recent approaches increasingly relying on machine learning. The current state-of-the-art, UCBShift (Zhou et al., 2023), combines sequence and structure alignment with a random forest model, using reference proteins and structural descriptors to guide predictions. Its latest iteration, UCBShift 2.0 (Ptaszek et al., 2024), extends this framework to include side-chain atoms alongside backbone predictions. Despite their effectiveness, these methods depend on reference-based similarity measures, which limits their generalizability.
Figure 4: Synthetic example showing the influence of a phenylalanine sidechain aromatic ring on surrounding chemical shifts. A. The magnitude of change of backbone CA chemical shifts over different ring orientations as predicted using the proposed MACE-based predictor (left) and UCBShift-X (right). A 77Å sphere indicates the radius from the ring center at which the influence of the ring current is expected to become negligible. Note that UCBShift-X predicts much longer-range, albeit small, ring influence extending beyond 2020Å. B. Locations of three nearby CA atoms; and C. their predicted chemical shifts vs. the ring orientation. Note the smooth $180^{\circ}$-periodic behavior of the MACE shift prediction and the decay of the effect scale with the distance from the ring.
@@
MLFF-based shift predictor. Chemical shift prediction is formulated as a transfer learning task over frozen MLFF embeddings, where a graph neural network takes the pretrained atomistic representations as input and predicts the chemical shift of a target atom. Separate models are trained for the N, CA, C, H, and HA backbone atoms, as well as the CB, CG, CD, CD2, CG2, and CE side-chain atoms. Dataset curation, model architecture, and training details are provided in Appendix E,  G.3.
@@
As evident from Fig. 3, the proposed predictor outperforms UCBShift2 for both backbone and side-chain heavy atoms, with the exception of the Hydrogen alpha atom. A comprehensive quantitative comparison of chemical shift prediction accuracy across different MLFF embeddings, ESM variants, and learned embeddings is provided in Tables B.5 and  B.6. We observe that MACE-based models consistently achieve the best performance.
@@
To further probe whether the shift predictor correctly captures local structural effects, we design three case studies: one is shown here, and the other two are provided in Appendix I for completeness.
@@
Effects of ring currents. An external magnetic field induces a flow of delocalized $\pi$ electrons in an aromatic ring, which in turn produces its own magnetic field. Because chemical shifts depend on the net magnetic field produced by the local structural environment, they are strongly influenced by such ring currents. As the phenylalanine ring has a C2 symmetry, states separated by rotation of the ring by $180^{\circ}$ are indistinguishable. To probe whether our shift predictor captures ring-current effects accurately, we synthetically modified the side chain of a Phe (residue $50$ in PDB ID: 1ZV6) by rotating its aromatic ring about the $\chi_{2}$ dihedral angle from $180^{\circ}$ to $180^{\circ}$. Fig. 4 depicts the studied environments and the predicted chemical shifts. Our MLFF-based shift predictor shows the expected $180^{\circ}$ periodicity (Haigh & Mallion, 1979) in the chemical shifts of backbone CA atoms in the vicinity of the ring, with the ring current influence decaying smoothly as the distance from the ring increases and becoming negligible after  77Å. In contrast, UCBShift extends this influence beyond the expected range and fails to reproduce the smoothness and periodicity expected in theory.
Figure 5: Lower likelihood environments result in larger chemical shift prediction error. Chemical shift prediction accuracy of CA (left) and N (right) atoms stratified by the KDE-estimated likelihood of the corresponding MACE descriptors. Depicted are the median and $25\%$-$50\%$ confidence intervals. Higher-likelihood environments correspond to lower prediction error and can be used as an uncertainty measure.
@@
Uncertainty estimation. We first build the unconditional likelihood model $p(\mathcal{X})$ for biomolecular environments on the training set used for shift prediction. We then evaluate this likelihood for every test environment and record the corresponding chemical shift prediction error. Fig. 5 reports the distribution of errors, stratified by likelihood. Lower likelihood environments yield higher errors, making likelihood a practical confidence measure representing, to the best of our knowledge, the first such score reported for chemical shift prediction.
@@
We summarize by concluding that MLFF-based shift prediction not only delivers better accuracy, but is also physics-grounded and allows us to provide confidence estimates for each prediction.
7 Interpreting MLFF representations